Determine the accelerations that result when a 20 N net force is applied to a 5 kg object and then to an 8 kg object.


1
Expert's answer
2022-02-22T08:52:36-0500

The Newton's second law says

"a_1=\frac{F}{m_1}=\frac{8\:\rm N}{5\:\rm kg}=1.6\:\rm m/s^2"

"a_2=\frac{F}{m_2}=\frac{8\:\rm N}{8\:\rm kg}=1\:\rm m/s^2"